OpenStudy (54tankcb):

Emily is playing a board game that has a spinner divided into equal sections numbered 1 to 18. The probability of the spinner landing on an even number or a multiple of 3 is

Mehek (mehek14):

How many multiples of 3 are there from 1-18?

OpenStudy (54tankcb):

6

Mehek (mehek14):

Good So we set up a fraction: \(\bf\dfrac{multiple~of~3}{total~numbers}=\dfrac{6}{18}=?\) Can you simplify that for me

OpenStudy (54tankcb):

\[\frac{ 1 }{ 6 }\]

Mehek (mehek14):

The denominator is wrong

OpenStudy (54tankcb):

oh my bad 3. 1 over 3

Mehek (mehek14):

Yes

Mehek (mehek14):

So the answer is \(\dfrac{1}{3}\)